OpenStudy (anonymous):

What is the slope of the line which passes through (1, –2) and (–1, 6)? 4 –4 2 –2

hero (hero):

Using y = mx + b, input both points into the formula separately to get -2 = m + b 6 = -m + b Combine both fractions together to get 4 = 2b Divide both sides by 2 to get 2 = b Substitute b back into the original equation: -2 = m + 2 subtract 2 from both sides to get: -2 - 2 = m -4 = m Since m represents slope, m = -4 means that the slope is -4

OpenStudy (anonymous):

-4